OpenStudy (anonymous):

please help Given the function f(x) = The quantity of 4x minus 2, divided by 3, which of the below expressions is correct? f−1(x) = The quantity of 2 minus 4x, divided by 3. f−1(x) = The quantity of 3x minus 2, divided by 4. f−1(x) = The quantity of negative 4x minus 2, divided by 3. f−1(x) = The quantity of 3x plus 2, divided by 4.

OpenStudy (anonymous):

f−1(x) = The quantity of 2 minus 4x, divided by 3. f−1(x) = The quantity of 3x minus 2, divided by 4. f−1(x) = The quantity of negative 4x minus 2, divided by 3. f−1(x) = The quantity of 3x plus 2, divided by 4.